# Modeling Platelet P2Y1/12 Pathway to Integrin Activation

### v1

@@ Abstract
Through experimental studies, many details of the pathway of integrin IIb3 activation by ADP during the platelet aggregation process have been mapped out. ADP binds to two separate G protein coupled receptors on platelet surfaces, leading to alterations in the regulation of the small GTPase RAP1. We seek to (1) gain insights into the relative contributions of both pathways to RAP1-mediated integrin activation and to (2) predict cell behavior in response to a continuous range of external agonist concentrations. To this end, we develop a dynamical systems model detailing the action of each protein in the two pathways up to the regulation of RAP1. We perform a parameter estimation using flow cytometry data to determine a number of unknown rate constants. We then validate with already published data; in particular, the model confirmed the effect of impaired P2Y1 receptor desensitization or reduced RASA3 expression on RAP1 activation. We then predict the effect of protein expression levels on integrin activation and show that components of the P2Y12 pathway are critical to the regulation of integrin. This model aids in our understanding of interindividual variability in platelet response to ADP and therapeutic P2Y12 inhibition. It also provides a more detailed view of platelet activation in the ongoing mathematical study of platelet aggregation.
@@
A detailed dynamical systems model of integrin IIb3 activation mediated by ADP is presented in this report. This model takes in years of experimental literature on the relevant pathways into simulations that are efficient to run and simple to manipulate parameters to match a platelet mutation. Unlike in experiments, our model can finely adjust any parameter and analyze how a platelet responds to any number of experimental settings. We present the utility of such a model through parameter study on the copy number of proteins key to integrin activation and show that integrin response changes more significantly in response to changes in the P2Y12 pathway over the P2Y1 pathway.
@@ Introduction
Several molecules serve as receptor ligands that lead to integrin activation (1). Most ligands, like thrombin or collagen, initiate pathways that leads to irreversible activation of the small protein RAP1, which complexes with the integrin on its cytosolic end and leads to a conformation change on the extracellular end. Signaling pathways for other receptor-ligand bindings have been extensively studied by both experimentalists and modelers. Dunster et al. modeled the activation signals associated with collagen binding the GPVI receptor (2) and Lenoci et al. and Sveshnikova et al. modeled the irreversible activation of small GAP protein RAP1 by the PAR1 receptor (3, 4).
ADP binds to two different receptors, the G-protein coupled receptors (GPCR) P2Y1 and P2Y12, and is unique in that it leads to a partial, transient activation of integrin that lasts only a few minutes. Experimental work has uncovered much about the pathways that lead to the activation of integrin. Activation of P2Y1 induces activation of RAP1 via the calcium-dependent RAP-GEF, CalDAG-GEFI (CDGI), while activation of P2Y12 induces RAP1 activation through the inhibition of the RAP-GAP, RASA3 (5). The regulation of RAP1 is vital to maintaining platelet quiescence in the absence of clot formation, and therefore understanding the individual contribution of CalDAG-GEFI and RASA3 in integrin activation is of interest. Other models have been developed to study platelet activation and aggregation mediated by ADP (6), but to the authors’ knowledge, no detailed signaling pathways for ADP dependent activation have been developed.
@@
In this work, we convert the collected information on both pathways to a system of ordinary differential equations (ODEs) that can be solved numerically to simulate a single platelet’s response to a given amount of P2Y1 and/or P2Y12 agonist. We use experimental data to estimate unknown parameters and validate the model against recently published data. We then explore how a simulated platelet responds to agonists under a change to the expression level of specified proteins. Unlike in experimental settings, we are able to tune expression with much finer precision to gain a more quantitative measure of sensitivity. We begin by describing the chemical signaling pathways upon platelet GPCR binding to ADP.
@@
The extracellular domains of the platelet P2Y1 and P2Y12 receptors bind to ADP in the blood plasma. Upon binding to ADP, these receptors act as GEF enzymes on the membrane-associated G proteins Gq and Gi, respectively. The G proteins then unbind from the receptor and bind to membrane-associated enzymes PLC and PI3K, respectively. Both PLC and PI3K use the lipid PIP2 as substrate: PLC converts PIP2 to the membrane-associated molecule DAG and the cytosolic molecule IP3, and PI3K converts PIP2 to membrane-associated PIP3. These inositol lipids are part of a large cycle of formation and degradation. IP3 binds to IP3 receptors (IP3R) embedded in the dense tubular system (DTS) membrane and triggers the release of calcium ions from the DTS into the cytosolic space. Platelets contain other intracellular calcium stores, e.g., the mitochondrial and acidic stores (7), which impact calcium levels on a significantly longer timescale and therefore are ignored for this study. The chemically-gated IP3R channel is a four-subunit protein; each subunit contains an activating binding site for IP3, an activating binding site for calcium, and a slower inactivating binding site for calcium (8). Thus, the initial release of calcium from the DTS leads to a stage of positive feedback followed by negative feedback.
At rest, cytosolic calcium levels are maintained by exchanges with the plasma, in which calcium is assumed to have a concentration around 1 . The model incorporates the action of the PMCA pump, a passive leak current on the plasma membrane, and the action of the SERCA pump across the DTS membrane.
Calcium’s effect on integrin activation is mediated through its binding to the GEF enzyme CalDAG-GEFI. CalDAG-GEFI contains two EF domains which each binds to calcium with a dissociation constant (9). Calcium-bound CalDAG-GEFI converts RAP1 from its inactive GDP-bound form to its active GTP-bound form. The active form of RAP1 complexes with a variety of proteins on the cytosolic tail of , namely TALIN and KINDLIN3, to trigger integrin activation. RASA3 is a membrane-associated protein constitutively active in converting active RAP1 to its inactive GDP-bound form. RASA3 can be inactivated by interacting with PIP3, although the exact mechanism has not been determined (10). CalDAG-GEFI and RASA3 have nearly equal copy numbers in mouse platelets (approximately 30,000 ) whereas RAP1 in its two major isoforms has a total copy number of approximately 200,000 (11).
@@
The model explicitly tracked the second messenger proteins PLC and PI3K in their inactive state and active (i.e., bound to G protein) state (15). Binding and unbinding of G proteins were tracked through Mass Action kinetics, with an additional irreversible hydrolysis term, in which the second messenger both converts the subunit to a GDP bound form and releases it into the cytosol. Once activated, PLC and PI3K act upon the phosphoinositol species PIP2, converting it to IP3 and PIP3, respectively. These inositol species are recycled through a variety of states by various other phosphatases and kinases (16, 17, 18, 19, 20, 21, 22, 23).
@@
The method begins by defining a hypercube in parameter space; for each dimension, the center is given by the average value of a specified protein’s copy number, and the half-width is given by the standard deviation of that protein’s copy number. Beginning at the center, the algorithm defines several paths to the boundary of the hypercube such that only one parameter changes with each step along a path. The paths were computed using Python’s SALib package (25, 26). Each vertex on a path specifies a parameter set used for a single simulation. In each simulation, we recorded the peak number of RAP1-GTP and then used finite differences to approximate the derivative of the maximum RAP1-GTP concentration with respect to the single parameter changed in that step. This process generated a sampling of the derivative of peak RAP1-GTP as a function of each parameter within the hypercube; for each parameter, we reported the mean and standard deviation of the approximation to the derivative. A large value of suggests a high sensitivity with respect to the given parameter, and a large value of suggests a high correlation between the given parameter and other parameters in the model. We again relied on SALib to compute and for each parameter.
@@
Figures 3b-c show a comparison between the fibrinogen binding data and the simulated RAP1-GTP spike using the lower bound approximation (see Methods) of the data in Figure 3b and the upper bound approximation of the data in Figure 3c. For each of these assumptions about the number of integrins activated by ADP, our estimation procedure sought to minimize the sum of the mean squared differences between data and simulation for both the WT and RASA3+/- cases. Our model approximated the data well when ADP was applied to the peak RAP1-GTP response. However, simulations showed a slower return to baseline than the experimental data. For estimated parameter values, see Supplemental Tables LABEL:tab:CaEstimatedParams and LABEL:tab:RAP1EstimatedParams.
@@
To gauge variations in platelet responses due to variations in agonist concentration, we performed computational experiments for a wide range of applied ADP and MRS2365 concentrations. In the corresponding physical experiments, the readout was the amount of the integrin probe JON/A bound irreversibly to activated integrin. As a proxy for the amount of JON/A bound at time we used the time-integral of the RAP1-GTP curve up to that time. Figure 8 shows the results of these experiments. As expected, for both agonists, the RAP1-GTP response increased as the concentration of the agonist was increased. This response was consistently greater for 340-0P platelets than for WT ones. This is attributed to the fact that for the P2Y platelets, the P2Y1 receptors remained active and continued to perform their GEF action on the Gq protein for the duration of the simulation. Consequently, the IP3, calcium, and RAP1-GTP responses are prolonged, and therefore, the area under the RAP1-GTP curve increased.
@@
We compared the results of Figures 2c and 4c in (27), recreated in Figure 8a and 8b, to the results of Figures 8c and 8d. First, comparing the experiments in which platelets are stimulated with ADP in Figure 8a and 8c, we saw an increase in the integrin response over the same range of applied ADP concentration for both wildtype and P2Y platelets. Additionally, we saw that for a given applied ADP concentration, the 340-0P platelet signal was consistently larger than the wildtype platelet. We note that in Figure 8b and 8d, we saw a similar relationship between mutant and wildtype platelets. Deviations between the model and experiment are evident when comparing simulated versus experimental mutant platelets stimulated with MRS2365, where the experiments show a more gradual increase in integrin activation as a function of applied agonist concentration compared to simulations. In general, comparing our model to results from (27) showed good agreement, even when a single experimental modification is considered. As more modifications are combined into simulated platelets, the results may deviate from experiments.
@@
The model allowed for the continuous variation of parameters, specifically protein copy numbers. We leveraged this to assess the effects of protein expression levels on the integrin activation response to a specified stimulus. In each set of simulations, we set the copy number of one of the proteins RASA3, CDGI, P2Y1, or P2Y12 to a value between 5% and 150% of its literature value (while holding other copy numbers at their literature values). We stimulated the platelet with 10 M ADP.
@@
Figure 9c plots the RAP1-GTP response as a function of a specified protein’s copy number, where we saw that the sensitivity of the response to RASA3 copy number variations was much greater than that for CDGI variations. Changes in the CDGI copy number led to a small and approximately linear change in RAP1 activation. In contrast, responses to decreases in the RASA3 copy number are nonlinear, with a sharp increase in sensitivity as the RASA3 copy number decreases to and beyond 50% of its baseline value.
Figure 10 shows a similar dichotomy between responses to variations in copy numbers of the ADP receptors P2Y1 and P2Y12. With P2Y12 at 150%, integrin activation was much greater at all times, and high levels were maintained for an extended period compared to 100%. With P2Y12 at 50%, integrin activation is weak. With P2Y1 at 50%, integrin activation was approximately half that of activation at 100%. There was only a small difference in integrin activation for P2Y1 100% and 150%. Examining model outputs, in this case, showed that although the amount of activated IP3 increased, the peak concentration of calcium did not increase significantly beyond 200 nM (data not shown).
@@ Conclusion
In this report, we developed a mathematical model of integrin activation on a platelet surface and fitted it to experimental data. Thanks to years of experimental studies, a chemical pathway has been defined sufficiently to allow such a model to exist. Previous platelet activation models rely on other major indicators, such as calcium signals or aggregometry readings (28, 29). While effective, inferring integrin activation information from only one of the two ADP-dependent pathways does not fully encapsulate the regulatory pathway’s complexity, making it difficult to extend to particular mutant platelets. Our model can be used in testing and making predictions without the need for a large number of experimental trials. It is quite economical to run; for reference, a single simulation takes approximately one minute to run ten simulated minutes on a standard personal computer. Thus, data generation is significantly faster than generation of an equivalent amount of data in the lab.
@@
In experiments where we varied the applied ADP concentration, we found that a platelet’s response saturates beyond 10 M, consistent with experiments. Our highly detailed model allowed us to examine the individual responses of each protein in time. We showed the usefulness of our model beyond directly comparing to experimental data by varying parameters that would be difficult to vary continuously in the lab, like protein expression levels, and making predictions on integrin activation. In particular, we showed that the total integrin response varies more significantly with changes to proteins in the P2Y12 pathway compared to the P2Y1 pathway.
This model can be used to conduct initial tests on the behavior of platelets under any form of experimental condition before running experiments. In addition, the outputs of this model describe integrin activation over time in far more detail than is typical in mathematical models for hemostasis and thrombosis (30, 31), where the tracking of platelet aggregation and fibrin formation is of interest. This may lead to improvements in these larger-scale models.

### v2

@@ Abstract
Through experimental studies, many details of the pathway of integrin activation by ADP during the platelet aggregation process have been mapped out. ADP binds to two separate G protein coupled receptors on platelet surfaces, leading to alterations in the regulation of the small GTPase RAP1. We seek to (1) gain insights into the relative contributions of both pathways to RAP1-mediated integrin activation and to (2) predict wildtype and mutated cell behavior in response to a continuous range of external agonist concentrations. To this end, we develop a dynamical systems model detailing the action of each protein in the two pathways up to the regulation of RAP1. We perform a parameter estimation using flow cytometry data to determine a number of unknown rate constants. We then validate with already published data; in particular, the model confirmed the effect of impaired P2Y1 receptor desensitization or reduced RASA3 expression on RAP1 activation. We then predict the effect of protein expression levels on integrin activation and show that components of the P2Y12 pathway are critical to the regulation of integrin. This model aids in our understanding of interindividual variability in platelet response to ADP and therapeutic P2Y12 inhibition. It also provides a more detailed view of platelet activation in the ongoing mathematical study of platelet aggregation.
@@
A detailed dynamical systems model of integrin activation mediated by ADP is presented. This model takes in years of experimental literature on the relevant pathways into simulations that are efficient to run and simple to manipulate parameters to match a platelet mutation. Our model can finely adjust any parameter and analyze how a platelet responds to any number of experimental settings. We present the model’s utility through parameter study on the expression levels of proteins key to integrin activation and show that integrin response changes more significantly in response to changes in the P2Y12 pathway over the P2Y1 pathway. This insight can be used to understand how certain antithrombotic therapies, like clopidogrel, can affect a platelet’s intracellular signaling pathways.
@@ Introduction
Several molecules serve as receptor ligands that lead to integrin activation (1). Most ligands, like thrombin or collagen, initiate pathways that lead to irreversible activation of the integrin. Signaling pathways for these receptor-ligand bindings have been studied by both experimentalists and modelers. Dunster et al. modeled the activation signals associated with collagen binding the GPVI receptor (2). Lenoci et al. (3) and Sveshnikova et al. (4) modeled platelet activation through the action of thrombin on the PAR1 receptor.
We focused our modeling efforts on platelet inside-out integrin activation induced by ADP, an autocrine/paracrine small molecule agonist released by activated platelets. Our main reasons for choosing ADP as the agonist include (i) its physiological relevance as a key recruiter of free-flowing platelets to a growing aggregate in vivo and is the target of certain antithrombotic therapies such as clopidogrel (5), (ii) the well-defined signaling pathways connecting the ADP receptors, P2Y1 and P2Y12, to integrin activation mediated by the adapter protein, Talin1, and (iii) a relatively weak and reversible integrin activation response that provides a perfect system to model both positive and negative modulation of these signaling pathways. Furthermore, we focus specifically on the effect of inside-out signaling in our first version of this model. We experimentally isolate the effect of inside-out signaling by using the fluorescent molecule JON/A(PE) (6), which selectively binds to the active conformation of the integrin and prevents outside-in signaling.
ADP binds to two platelet GPCRs, the Gq-coupled P2Y1 and Gi-coupled P2Y12 receptors. Activation of P2Y1 induces activation of phospholipase C (PLC) and formation of Ca2+ and diacylglycerol (DAG), while activation of P2Y12 induces PI3 kinase-dependent formation phosphatidylinositol-3,4,5-trisphosphate (PIP3). Ca2+ activates the RAP-GEF, CalDAG-GEFI (CDGI), while PIP3 inhibits the RAP-GAP, RASA3 (7). Activation of both pathways, but not P2Y1 or P2Y12 signaling alone, leads to robust and prolonged activation of small GTPase RAP1, a key regulatory component of the integrin activation complex (IAC) (8, 9, 10). The IAC consists of RAP1 and adapter proteins such as Talin1 (11). Following RAP1 activation, these proteins assemble at the intracellular tail of the integrin to induce a conformational change required for ligand binding. Once RAP1-GTP is converted to its GDP bound form, however, these proteins quickly unbind, and the integrin returns to its inactive conformation. Thus, the dynamics of RAP1 activation can be used as a proxy for integrin activation. It is also important to note that ADP-dependent mobilization of intracellular calcium stores is also linked to the influx of calcium from the extracellular space. However, this so-called store-operated calcium entry (SOCE) does not impact ADP-induced integrin inside-out activation (12). Furthermore, P2Y12 signaling leads to decreased levels of cyclic AMP, a second messenger that leads to the inhibition of various platelet activation responses (13). We have not included these pathways in our model as our previous work demonstrated that signals directly affecting CalDAG-GEFI and RASA3 function are most relevant for RAP1-dependent integrin inside-out activation.
@@
Other models (14) have been developed to study integrin activation and aggregation mediated by ADP, but to the authors’ knowledge, no detailed models of signaling pathways for ADP-dependent activation have been developed. In this work, we convert the collected information on both pathways to a system of ordinary differential equations (ODEs) that can be solved numerically to simulate a single platelet’s response to a given amount of P2Y1 and/or P2Y12 agonist. We use experimental data to estimate unknown parameters and validate the model against recently published data. We then explore how a simulated platelet responds to agonists under a change to the expression level of P2Y1, P2Y12, RASA3, and CalDAG-GEFI. Unlike in experimental settings, we are able to tune expression with much finer precision to gain a more quantitative measure of sensitivity. We begin by describing the chemical signaling pathways triggered upon platelet GPCR binding to ADP.
@@
The extracellular domains of the platelet P2Y1 and P2Y12 receptors bind to ADP in the blood plasma. Upon binding to ADP, these receptors act as GEF enzymes on the membrane-associated G proteins Gq and Gi, respectively. The G proteins then unbind from the receptor and bind to membrane-associated enzymes PLC and PI3K, respectively. Both PLC and PI3K use the lipid PIP2 as substrate: PLC converts PIP2 to the membrane-associated molecule DAG and the cytosolic molecule IP3, and PI3K converts PIP2 to membrane-associated PIP3. These inositol lipids are part of a large cycle of formation and degradation. IP3 binds to IP3 receptors (IP3R) embedded in the dense tubular system (DTS) membrane and triggers the release of calcium ions from the DTS into the cytosolic space.
The chemically-gated IP3R channel is a four-subunit protein; each subunit contains an activating binding site for IP3, an activating binding site for calcium, and a slower inactivating binding site for calcium (15). Thus, the initial release of calcium from the DTS leads to a stage of positive feedback followed by negative feedback. In resting platelets, the cytosolic Ca2+ concentration is 20-100 nM, a level that is maintained by active sequestration of Ca2+ in the DTS and the extracellular space. In addition to the IP3R channel, the model incorporates the active PMCA pump and passive leak current on the plasma membrane, and the active SERCA pump across the DTS membrane.
Calcium’s effect on integrin activation is mediated through its binding to the GEF enzyme CalDAG-GEFI. CalDAG-GEFI contains two EF domains, which each binds calcium with a dissociation constant (16). Calcium-bound CalDAG-GEFI converts RAP1 from its inactive GDP-bound form to its active GTP-bound form. The active form of RAP1 complexes with a variety of proteins on the cytosolic tail of , for example, TALIN, to trigger integrin activation. RASA3 is a membrane-associated protein constitutively active in converting active RAP1 to its inactive GDP-bound form. RASA3 can be inactivated by interacting with PIP3, although the exact mechanism has not been determined (17). CalDAG-GEFI and RASA3 have nearly equal copy numbers in mouse platelets (approximately 30,000 ), whereas RAP1 in its two major isoforms has a total copy number of approximately 200,000 (18).
@@
The model explicitly tracked the second messenger proteins PLC and PI3K in their inactive state and active (i.e., bound to G protein) state (22). Binding and unbinding of G proteins were tracked through Mass Action kinetics, with an additional irreversible hydrolysis term, in which the second messenger both converts the subunit to a GDP bound form and releases it into the cytosol. Once activated, PLC and PI3K act upon the phosphoinositol species PIP2, converting it to IP3 and PIP3, respectively. These inositol species are recycled through a variety of states by various other phosphatases and kinases (23, 24, 25, 26, 27, 28, 29, 30).
@@
The method begins by defining a hypercube in parameter space; for each dimension, the center is given by the average value of a specified protein’s copy number, and the half-width is given by the standard deviation of that protein’s copy number. Beginning at the center, the algorithm defines several paths to the boundary of the hypercube such that only one parameter changes with each step along a path. The paths were computed using Python’s SALib package (32, 33). Each vertex on a path specifies a parameter set used for a single simulation. In each simulation, we recorded the peak number of RAP1-GTP and then used finite differences to approximate the derivative of the maximum RAP1-GTP concentration with respect to the single parameter changed in that step. This process generated a sampling of the derivative of peak RAP1-GTP as a function of each parameter within the hypercube; for each parameter, we reported the mean and standard deviation of the approximation to the derivative. A large value of suggests a high sensitivity with respect to the given parameter, and a large value of suggests a high correlation between the given parameter and other parameters in the model. We again relied on SALib to compute and for each parameter.
@@
Figures 3b-c show a comparison between the fibrinogen binding data and the simulated RAP1-GTP spike using the lower bound approximation (see Methods) of the data in Figure 3b and the upper bound approximation of the data in Figure 3c. For each of these assumptions about the number of integrins activated by ADP, our estimation procedure sought to minimize the sum of the mean squared differences between data and simulation for both the WT and RASA3+/- cases. Our model approximated the data well when ADP was applied to the peak RAP1-GTP response. However, simulations showed a slower return to baseline than the experimental data. For estimated parameter values, see Supplemental Tables S1 and S2.
@@
To gauge variations in platelet responses due to variations in agonist concentration, we performed computational experiments for a wide range of applied ADP and MRS2365 concentrations. In the corresponding physical experiments, the readout was the amount of the integrin probe JON/A bound irreversibly to activated integrin. As a proxy for the amount of JON/A bound at time , we used the time-integral of the RAP1-GTP curve up to that time. Figure 8 shows the results of these experiments. As expected, for both agonists, the RAP1-GTP response increased as the concentration of the agonist was increased. This response was consistently greater for 340-0P platelets than for WT ones. This is attributed to the fact that for the P2Y platelets, the P2Y1 receptors remained active and continued to perform their GEF action on the Gq protein for the duration of the simulation. Consequently, the IP3, calcium, and RAP1-GTP responses are prolonged, and therefore, the area under the RAP1-GTP curve increased.
@@
We compared the results of Figures 2c and 4c in (10), recreated in Figure 8a and 8b, to the results of Figures 8c and 8d. First, comparing the experiments in which platelets are stimulated with ADP in Figure 8a and 8c, we saw an increase in the integrin response over the same range of applied ADP concentration for both wildtype and P2Y platelets. Additionally, we saw that for a given applied ADP concentration, the 340-0P platelet signal was consistently larger than the wildtype platelet. We note that in Figure 8b and 8d, we saw a similar relationship between mutant and wildtype platelets. Deviations between the model and experiment are evident when comparing simulated versus experimental mutant platelets stimulated with MRS2365, where the experiments show a more gradual increase in integrin activation as a function of applied agonist concentration compared to simulations. Comparing our model to results from (10) showed good agreement, even when a single experimental modification is considered. As more modifications are combined into simulated platelets, the results may deviate from experiments.
@@
The model allowed for the continuous variation of parameters, specifically protein copy numbers. We leveraged this to assess the effects of protein expression levels on the integrin activation response over a range of ADP stimuli. In each set of simulations, we set the copy number of one of the proteins RASA3, CDGI, P2Y1, or P2Y12 to a value between 5% and 150% of its literature value (while holding other copy numbers at their literature values).
@@
Figure 9c plots the RAP1-GTP response of platelets with specified expression level changes as a function of ADP concentration. We found that as we vary the expression of RASA3, the dose-response curve deviates further from the WT curve than for equivalent variations in the expression of CDGI. Furthermore, the variation in dose-response occurs across applied ADP concentrations. For example, at 10 nM ADP, there are modest increases in the integrin response as we vary expression levels within 50% of baseline and much more dramatic increases as we decrease RASA3 expression down to 5%. This result mirrors behaviors seen experimentally in Stefanini et al. (7), where a specific mouse platelet mutation resulting in 5% RASA3 expression led to an increase in both the baseline and stimulated integrin response. When we examined the integrin response as a function of expression level, we found that the dose-response curve scales approximately linearly with changes in CDGI expression (Supplemental Figure S1b). In contrast, platelet response as a function of RASA3 expression is highly nonlinear, with a dramatic dose-response curve as the RASA3 copy number decreases to and below 50% of its baseline value.
We next were interested in determining how platelets would respond to ADP when their receptor expression levels varied. In Figure 10, we vary the expression of P2Y1 and P2Y12 individually to see the platelet’s response to ADP over time and the total dose response. We expected that the integrin response should be positively correlated with the expression levels of either receptor, and Figures 10a-b agree. Examining model outputs further verifies that the calcium timecourse is affected by variation in P2Y1 expression levels but not by variation in P2Y12 expression levels (data not shown). Figure 10c shows the dose-response curves for different receptor expression levels. Similar to RASA3 in Figure S1b, we found that variation in P2Y12 yielded larger changes to the dose-response curve than in P2Y1.
Interestingly, Figures 10b-c show that an increase in the P2Y1 receptor to 150% leads to an almost negligible change in integrin response. Further analysis of model outputs shows that although increasing P2Y1 expression led to an increase in peak IP3 concentrations, the peak concentration of calcium did not increase significantly beyond 200 nM (data not shown). We show in Supplemental Figure S2a that the relative integrin response between 100% and 150% P2Y1 expression is more significant when 0.5 M ADP is used due to a difference in peak cytosolic calcium levels. Similar results were seen experimentally in Heckler et al. (34), in which they overexpress P2Y1 by 84% and notice an increase in aggregation response to low doses of ADP.
@@ Conclusion
In this report, we developed a mathematical model of inside-out integrin activation on a platelet surface and fitted it to experimental data. Thanks to years of experimental studies (1), a chemical pathway has been defined sufficiently to allow such a model to exist. Previous platelet activation models rely on other major indicators, such as calcium signals or aggregometry readings (35, 36). While effective, inferring integrin activation information from only one of the two ADP-dependent pathways does not fully encapsulate the regulatory pathway’s complexity, making it difficult to extend to particular mutant platelets. Our model can be used in testing and making predictions without the need for a large number of experimental trials. It is quite economical to run; for reference, a single simulation takes approximately one minute to run ten simulated minutes on a standard personal computer. Thus, data generation is significantly faster than generation of an equivalent amount of data in the lab.
The results in this work rely on a combination of kinetic rates found in the literature or estimated from data. Copy numbers of each protein were based on proteomics analysis of mouse platelets (18). In some cases, due to a lack of available data, rate constants associated with a protein isoform found in human platelets were used (23, 25, 26, 28, 30). This choice allows us to estimate fewer parameters under the assumption that the action of the protein isoform found in mouse platelets acts similarly on its substrate compared to these human platelet isoforms.
@@
In experiments where we varied the applied ADP concentration, we found that a platelet’s response saturates beyond 10 M, consistent with experiments. Our highly detailed model allowed us to examine the individual responses of each protein in time. We showed the usefulness of our model beyond directly comparing to experimental data by varying parameters that would be difficult to vary continuously in the lab, like protein expression levels, and making predictions on integrin activation. In particular, we showed that the total integrin response varies more significantly with changes to proteins in the P2Y12 pathway compared to the P2Y1 pathway. By decreasing the expression level of RASA3 to well below 50%, we found that both a resting and stimulated platelet presented a marked increase in integrin response, similar to results seen in Stefanini et al. (7). By changing expression levels of the receptors, we show that the integrin response can be modulated only above 0.1 M ADP. In Supplemental Figure S2a, we show that when using 50 nM ADP, we start to see a change in the integrin response between 100% and 150% expression of P2Y1, akin to the results of Hechler et al (34) in which they overexpress P2Y1.
The results from this study can be used to understand how certain pharmacological agents affect the integrin activation process. In particular, the P2Y12 antagonist clopidogrel is taken long-term as an antithrombotic treatment, and its action is analogous to a reduction in the overall expression of P2Y12. More detailed modeling is needed to show this connection explicitly, and will be the subject of future work.
This model can be used to conduct initial tests on the behavior of platelets under any form of experimental condition before running experiments. In addition, the outputs of this model describe integrin activation over time in far more detail than is typical in mathematical models for hemostasis and thrombosis (37, 38), where the tracking of platelet aggregation and fibrin formation is of interest. This may lead to improvements in these larger-scale models.
@@
Declaration of Interests
The authors declare no competing interests.